Can u pease help me w/ the following problem... 4(3x-4)=8(17-8)
Answer by checkley71(8403)

4(3X-4)=8(17-8)
12X-16=8*9
12X-16=72
12X=72+16
12X=88
X=88/12
X=22/3
PROOF
4(3*22/3-4)=8(17-8)
4(22-4)=8(9)
4*18=72
72=72